1 − sq. root 3i. Write the trigonometric form of the complex number. (Let 0 ≤ θ < 2π.)
Answer #1

AND - = I-Bi 2= i-ins Grillen complex Nomber To Wanite the neember in trigonometric from he meeed 2 = atib from 42 ar and o a